OpenStudy (anonymous):

Factor completely 3x3 + 6x2 - 24x. (4 points) 3x(x - 4)(x + 2) 3x(x + 4)(x - 2) 3x(x - 3)(x + 2) 3x(x + 3)(x + 2)

OpenStudy (solomonzelman):

first take out the 'x' what do you get after doing so>?

OpenStudy (anonymous):

3+6-24

OpenStudy (compassionate):

Eh, what he's saying is factor x throughout the expression via LCD/M

OpenStudy (compassionate):

\[3x^3 + 6x^2 - 24x = x(3x^2 + 5x - 24)\] Now you're simply left with, on the right side, a binomial

OpenStudy (compassionate):

\[3x^2 + 6x - 24\]

OpenStudy (anonymous):

ok.. whats the next step?

OpenStudy (compassionate):

\[x(3x^2 + 6 - 24) = x(3x \pm ?)(x \pm ?)\] You are simply solving it as if it were a binomial.

OpenStudy (anonymous):

ok. let me write it out in my note book

Nnesha (nnesha):

there is a typo \[x( 3x^2 + 6x -24)\] suppose to be this

OpenStudy (anonymous):

._. ugh

Nnesha (nnesha):

yep first step is to find GCF ( greatest common factor ) like they told you that x is one of common factor so basically you have to divide whole equation by x